$SPAD/src/input ico.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool ico.output
)set message test on
)set message auto off
)clear all
--S 1 of 65
)se exp add con InnerTrigonometricManipulations
--R
--R   InnerTrigonometricManipulations is now explicitly exposed in frame
--R      initial
--E 1
--S 2 of 65
exp(%i*2*%pi/5)
--R
--R
--R          2%i %pi
--R          -------
--R             5
--R   (1)  %e
--R                                           Type: Expression(Complex(Integer))
--E 2
--S 3 of 65
FG2F %
--R
--R
--R               +---+
--R          2%pi\|- 1
--R          ----------
--R               5
--R   (2)  %e
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 65
% -1
--R
--R

--R               +---+
--R          2%pi\|- 1
--R          ----------
--R               5
--R   (3)  %e           - 1
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 65
complexForm %
--R
--R
--R            2%pi            2%pi
--R   (4)  cos(----) - 1 + sin(----)%i
--R              5               5
--R                                           Type: Complex(Expression(Integer))
--E 5
--S 6 of 65
norm %
--R
--R
--R            2%pi 2       2%pi 2        2%pi
--R   (5)  sin(----)  + cos(----)  - 2cos(----) + 1
--R              5            5             5
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 65
simplify %
--R
--R
--R               2%pi
--R   (6)  - 2cos(----) + 2
--R                 5
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 65
s:=sqrt %
--R
--R
--R         +----------------+
--R         |       2%pi
--R   (7)   |- 2cos(----) + 2
--R        \|         5
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 65

ph:=exp(%i*2*%pi/5)
--R
--R
--R          2%i %pi
--R          -------
--R             5
--R   (8)  %e
--R                                           Type: Expression(Complex(Integer))
--E 9
--S 10 of 65
A1:=complex(1,0)
--R
--R
--R   (9)  1
--R                                                       Type: Complex(Integer)
--E 10
--S 11 of 65
A2:=A1*ph
--R
--R
--R           2%i %pi
--R           -------
--R              5
--R   (10)  %e
--R                                           Type: Expression(Complex(Integer))
--E 11
--S 12 of 65
A3:=A2*ph
--R
--R
--R            2%i %pi 2
--R            -------
--R               5
--R   (11)  (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 12
--S 13 of 65
A4:=A3*ph
--R
--R
--R            2%i %pi 3
--R            -------
--R               5
--R   (12)  (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 13

--S 14 of 65
A5:=A4*ph
--R
--R
--R            2%i %pi 4
--R            -------
--R               5
--R   (13)  (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 14
--S 15 of 65
ca1:=map(numeric , complexForm FG2F simplify A1)
--R
--R
--R   (14)  1.0
--R                                                         Type: Complex(Float)
--E 15
--S 16 of 65
ca2:=map(numeric , complexForm FG2F simplify A2)
--R
--R
--R   (15)  0.3090169943 749474241 + 0.9510565162 9515357212 %i
--R                                                         Type: Complex(Float)
--E 16
--S 17 of 65
ca3:=map(numeric ,complexForm FG2F simplify A3)
--R
--R
--R   (16)  - 0.8090169943 749474241 + 0.5877852522 9247312917 %i
--R                                                         Type: Complex(Float)
--E 17
--S 18 of 65
ca4:=map(numeric ,complexForm FG2F simplify A4)
--R
--R
--R   (17)  - 0.8090169943 7494742411 - 0.5877852522 9247312917 %i
--R                                                         Type: Complex(Float)
--E 18
--S 19 of 65
ca5:=map(numeric ,complexForm FG2F simplify A5)
--R
--R
--R   (18)  0.3090169943 749474241 - 0.9510565162 9515357212 %i
--R                                                         Type: Complex(Float)

--E 19
--S 20 of 65
B1:=A1*exp(2*%i*%pi/10)
--R
--R
--R           %i %pi
--R           ------
--R              5
--R   (19)  %e
--R                                           Type: Expression(Complex(Integer))
--E 20
--S 21 of 65
B2:=B1*ph
--R
--R
--R           %i %pi  2%i %pi
--R           ------  -------
--R              5       5
--R   (20)  %e      %e
--R                                           Type: Expression(Complex(Integer))
--E 21
--S 22 of 65
B3:=B2*ph
--R
--R
--R           %i %pi   2%i %pi 2
--R           ------   -------
--R              5        5
--R   (21)  %e      (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 22
--S 23 of 65
B4:=B3*ph
--R
--R
--R           %i %pi   2%i %pi 3
--R           ------   -------
--R              5        5
--R   (22)  %e      (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 23
--S 24 of 65
B5:=B4*ph
--R
--R

--R           %i %pi   2%i %pi 4
--R           ------   -------
--R              5        5
--R   (23)  %e      (%e       )
--R                                           Type: Expression(Complex(Integer))
--E 24
--S 25 of 65
cb1:=map (numeric ,complexForm FG2F simplify B1)
--R
--R
--R   (24)  0.8090169943 749474241 + 0.5877852522 9247312917 %i
--R                                                         Type: Complex(Float)
--E 25
--S 26 of 65
cb2:=map (numeric ,complexForm FG2F simplify B2)
--R
--R
--R   (25)  - 0.3090169943 749474241 + 0.9510565162 9515357212 %i
--R                                                         Type: Complex(Float)
--E 26
--S 27 of 65
cb3:=map (numeric ,complexForm FG2F simplify B3)
--R
--R
--R   (26)  - 1.0
--R                                                         Type: Complex(Float)
--E 27
--S 28 of 65
cb4:=map (numeric ,complexForm FG2F simplify B4)
--R
--R
--R   (27)  - 0.3090169943 7494742409 - 0.9510565162 9515357212 %i
--R                                                         Type: Complex(Float)
--E 28
--S 29 of 65
cb5:=map (numeric ,complexForm FG2F simplify B5)
--R
--R
--R   (28)  0.8090169943 7494742411 - 0.5877852522 9247312916 %i
--R                                                         Type: Complex(Float)
--E 29
--S 30 of 65
u:=numeric sqrt(s*s-1)
--R

--R
--R   (29)  0.6180339887 4989484821
--R                                                                  Type: Float
--E 30
--S 31 of 65
p0:=point([0,0,u+1/2])@Point(SF)
--R
--R
--R   (30)  [0.,0.,1.1180339887498949]
--R                                                     Type: Point(DoubleFloat)
--E 31
--S 32 of 65
p1:=point([real ca1,imag ca1,0.5])@Point(SF)
--R
--R
--R   (31)  [1.,0.,0.5]
--R                                                     Type: Point(DoubleFloat)
--E 32
--S 33 of 65
p2:=point([real ca2,imag ca2,0.5])@Point(SF)
--R
--R
--R   (32)  [0.30901699437494745,0.95105651629515353,0.5]
--R                                                     Type: Point(DoubleFloat)
--E 33
--S 34 of 65
p2:=point([real ca2,imag ca2,0.5])@Point(SF)
--R
--R
--R   (33)  [0.30901699437494745,0.95105651629515353,0.5]
--R                                                     Type: Point(DoubleFloat)
--E 34
--S 35 of 65
p3:=point([real ca3,imag ca3,0.5])@Point(SF)
--R
--R
--R   (34)  [- 0.80901699437494734,0.58778525229247314,0.5]
--R                                                     Type: Point(DoubleFloat)
--E 35
--S 36 of 65
p4:=point([real ca4,imag ca4,0.5])@Point(SF)
--R
--R
--R   (35)  [- 0.80901699437494734,- 0.58778525229247303,0.5]

--R                                                     Type: Point(DoubleFloat)
--E 36
--S 37 of 65
p5:=point([real ca5,imag ca5,0.5])@Point(SF)
--R
--R
--R   (36)  [0.30901699437494745,- 0.95105651629515364,0.5]
--R                                                     Type: Point(DoubleFloat)
--E 37
--S 38 of 65
p6:=point([real cb1,imag cb1,-0.5])@Point(SF)
--R
--R
--R   (37)  [0.80901699437494745,0.58778525229247314,- 0.5]
--R                                                     Type: Point(DoubleFloat)
--E 38
--S 39 of 65
p7:=point([real cb2,imag cb2,-0.5])@Point(SF)
--R
--R
--R   (38)  [- 0.3090169943749474,0.95105651629515353,- 0.5]
--R                                                     Type: Point(DoubleFloat)
--E 39
--S 40 of 65
p8:=point([real cb3,imag cb3,-0.5])@Point(SF)
--R
--R
--R   (39)  [- 1.,0.,- 0.5]
--R                                                     Type: Point(DoubleFloat)
--E 40
--S 41 of 65
p9:=point([real cb4,imag cb4,-0.5])@Point(SF)
--R
--R
--R   (40)  [- 0.3090169943749474,- 0.95105651629515364,- 0.5]
--R                                                     Type: Point(DoubleFloat)
--E 41
--S 42 of 65
p10:=point([real cb5,imag cb5,-0.5])@Point(SF)
--R
--R
--R   (41)  [0.80901699437494745,- 0.58778525229247303,- 0.5]
--R                                                     Type: Point(DoubleFloat)
--E 42

--S 43 of 65
p11:=point([0,0,-u-1/2])@Point(SF)
--R
--R
--R   (42)  [0.,0.,- 1.1180339887498947]
--R                                                     Type: Point(DoubleFloat)
--E 43
--S 44 of 65
space:=create3Space()$ThreeSpace DFLOAT
--R
--R
--R   (43)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 44
--S 45 of 65
polygon(space,[p0,p1,p2])
--R
--R
--R   (44)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 45
--S 46 of 65
polygon(space,[p0,p2,p3])
--R
--R
--R   (45)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 46
--S 47 of 65
polygon(space,[p0,p3,p4])
--R
--R
--R   (46)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 47
--S 48 of 65
polygon(space,[p0,p4,p5])
--R
--R
--R   (47)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 48
--S 49 of 65

polygon(space,[p0,p5,p1])
--R
--R
--R   (48)  3-Space with 5 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 49
--S 50 of 65
polygon(space,[p1,p6,p2])
--R
--R
--R   (49)  3-Space with 6 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 50
--S 51 of 65
polygon(space,[p2,p7,p3])
--R
--R
--R   (50)  3-Space with 7 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 51
--S 52 of 65
polygon(space,[p3,p8,p4])
--R
--R
--R   (51)  3-Space with 8 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 52
--S 53 of 65
polygon(space,[p4,p9,p5])
--R
--R
--R   (52)  3-Space with 9 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 53
--S 54 of 65
polygon(space,[p5,p10,p1])
--R
--R
--R   (53)  3-Space with 10 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 54
--S 55 of 65
polygon(space,[p2,p6,p7])
--R

--R
--R   (54)  3-Space with 11 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 55
--S 56 of 65
polygon(space,[p3,p7,p8])
--R
--R
--R   (55)  3-Space with 12 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 56
--S 57 of 65
polygon(space,[p4,p8,p9])
--R
--R
--R   (56)  3-Space with 13 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 57
--S 58 of 65
polygon(space,[p5,p9,p10])
--R
--R
--R   (57)  3-Space with 14 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 58
--S 59 of 65
polygon(space,[p1,p10,p6])
--R
--R
--R   (58)  3-Space with 15 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 59
--S 60 of 65
polygon(space,[p6,p11,p7])
--R
--R
--R   (59)  3-Space with 16 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 60
--S 61 of 65
polygon(space,[p7,p11,p8])
--R
--R
--R   (60)  3-Space with 17 components

--R                                                Type: ThreeSpace(DoubleFloat)
--E 61
--S 62 of 65
polygon(space,[p8,p11,p9])
--R
--R
--R   (61)  3-Space with 18 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 62
--S 63 of 65
polygon(space,[p9,p11,p10])
--R
--R
--R   (62)  3-Space with 19 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 63
--S 64 of 65
polygon(space,[p10,p11,p6])
--R
--R
--R   (63)  3-Space with 20 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 64
--S 65 of 65
makeViewport3D(space,title=="Icosahedron")
--R
--R   Transmitting data...
--R
--R   (64)  ThreeDimensionalViewport: "Icosahedron"
--R                                               Type: ThreeDimensionalViewport
--E 65
)spool
)lisp (bye)

———-

References^{[1] nothing}

15